OpenStudy (anonymous):

Simplify 32 to the negative 3 fifths power

OpenStudy (anonymous):



OpenStudy (campbell_st):

32 = 2^5 so using index notation (2^5)^-3/5 multiply the powers gives 2^-3 or 1/8

OpenStudy (anonymous):

Eat the elephant one bite at a time:\[32^{-3/5} =(\sqrt[5]{32})^{-3}=(\sqrt[5]{2^5})^{-3}=2^{-3}=1/8\]

OpenStudy (anonymous):

Campbell, I think your method is probably more elegant. Nice work.